OpenStudy (anonymous):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21.

hartnn (hartnn):

y = 6(x+2) − 3 can you isolate x from that ? start by adding 3 to both sides

OpenStudy (shamil98):

First, distribute the 6(x+2) -3 6x + 12 -3 f(x) = 6x + 9 the inverse of this would be f(x) = 6x - 9 i think.

hartnn (hartnn):

another way id to put 6(x+2) − 3 = 21 and find 'x' but i would suggest you to first find the inverse function

OpenStudy (shamil98):

wait actually the inverse is f^-1(x) = 1/6 x - 9 right?.. my alg 2 is rusty..

hartnn (hartnn):

nopes, incorrect....

OpenStudy (anonymous):

I got 2 which was right.. Thanks!

hartnn (hartnn):

yes, 2 is correct final answer.

OpenStudy (shamil98):

the inverse function is , x - 9 / 6 21 - 9 = 12/6 = 2

OpenStudy (jdoe0001):

y == 6x + 9 inverse would be x= 6y + 9